The future value of a​ $10,000 annuity deposited at 12 percent compounded annually for each of next 5 years​ is: ​ (Round to the nearest whole​ dollar)

A.

​$40,376    

B.

​$63,528

C.

​$71,154   

D.

​$36,050   

The future value of an ordinary annuity of​ $1,000 each year for 10​ years, deposited at 3​ percent, is:​ (Round to the nearest whole​ dollar)

A.

​$8,530   

B.

​$11,464   

C.

​$11,808    

D.

​$10,000

Peter will receive​ $1,200 at the beginning of each of the next seven years. What is the future value of this​ annuity, assuming the interest rate is​ 9% compounded​ annually? (Round to the nearest whole​ dollar)

A.

​$8,400

B.

​$13,234   

C.

​$1,200   

D.

​$12,034   

E.

​$11,041    
Answer #1

Question 1:

[(1+г) — 1 FV of Annuity %3D Р т Р- Рeriodic Pаутent г3 rate per period п 3 питber o ff periods

(10.12) 1 FV 10000 0.12

FV = 10000 * 6.3528

FV = $63,528 (Option B)

Question 2:

Using the same formula as above,

(10.03)10 1 FV 1000 * 0.03

FV = 1000 * 11.4639

FV = $11,464 (Option B)

Question 3:

For FV of annuity due

[(1+r)- 1 FV of Annuity Due = (1 +r) x P P = Periodic Payment rrate per period n =number of periods

(10.09) 1 FV (10.09) 1200 0.09

FV = 1.09 * 1200 * 9.2004

FV = $12,034 (Option D)